Question 1. (15 points) Suppose that Ken cares only about bathing suits (B) and flip-flops (F). His utility function is U = B 0.6F 0.4. The price of bathing suits are $10, and the price of flip-flops are $4. Ken has a budget of $100. (For Version 2, Income is $200 but prices remain the same).

(a) (4 points) Draw a graph containing Ken’s budget line with bathing suits (B) on the x-axis and flip-flops (F) on the y-axis. Define the slope of this line.

(b) (2 points) Define the two conditions that need to be met for a consumption allocation {B ∗ , F ∗} to maximization consumption.

(c) (4 points) What is the marginal utility of good B (MUB) and of good F (MUF) for Ken. 2

(d) (5 points) What is the optimal choice of B and F for Ken to maximize his utility?
Answer #1

AND U=Borbon budget line 100 = 10B + 4 F Version 2 - 200 = 10 B+ 4F al m 100210 B+U A on up = 100-10B an f = 100 - 10 = 25-2.

100= 10 B+ 4 . 5 B an 303+20B 100 on 3 B = 100x3 = 6 and f = 5 x6 =10